OpenStudy (anonymous):

Use the Binomial Theorem to expand the expression (2y-3x)^5

OpenStudy (anonymous):

first find the fifth level of pascal's triangle for the coefficients

OpenStudy (anonymous):

do you know it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lol google!

OpenStudy (anonymous):

there is tons even "images" will be good the fifth level is the one that starts 1, 5

OpenStudy (anonymous):

did you find it yet?

OpenStudy (blurbendy):

binomial theorem of (a + b)^n = \[\sum_{k=0}^{n} \left(\begin{matrix}n \\ k\end{matrix}\right) a^{n-k}b^{k}\]

OpenStudy (anonymous):

http://www.mathsisfun.com/pascals-triangle.html

OpenStudy (anonymous):

the fifth level is the one with the number 1 , 5, 10, 10, 5, 1

OpenStudy (anonymous):

ok i got it, thanks again

OpenStudy (anonymous):

those are you coefficients then the powers have to add up to 5, so you can start with \[ (2y-3x)^5\] \[=(2y)^5-5(2y)^4(3x)+10(2y)^3(3x)^2-10(2y)^2(3x)^3+5(2y)(3x)^4-(3y)^5\]

OpenStudy (anonymous):

then you have a bunch of multiplication to do if it was me, i would cheat

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=+%282y-3x%29^5